The next five questions refer to the following situation: Consider two coun- tries, Westeros and Meereen. They each produce two goods: slaves, and wine. In one week, Westeros can either capture 2,000 slaves, or produce 3,000 kg of wine. Also, in one week, Meereen can produce either 3,000 slaves, or produce 1,000 kg of wine.1 1. Which country has absolute advantage in slaves? (A) Meereen (B) Westeros (C) Neither (D) Both 2. Which country has absolute advantage in wine? (A) Meereen B) Westeros (C) Neither (D) Both3. Which country has comparative advantage in slaves? (A) Meereen (B) Westeros (C) Neither (D) Both 4. If each country specializes in the good for which it has comparative advantage, how much will each produce? (A) Westeros will produce 2,000 slaves, and Meereen will produce 1,000 kg of wine (B) Westeros will produce 2,000 slaves, and Meereen will produce 3,000 slaves. (C) Westeros will produce 3,000 kg of wine, and Meereen will produce 1,000 kg of wine. (D) Westeros will produce 3,000 kg of wine, and Meereen will produce 3,000 slaves 5. Suppose, initially, that there is no trade between Westeros and Meereen. Westeros produces 1,000 slaves, and makes 1,500 kg of wine. Meereen produces 1,500 slaves, and 500 kg of wine. Now suppose that each country specializes in the good for which it has comparative advantage. The two countries then decide to trade with each other. If one slave can be exchanged for one kg of wine, can each country consume more goods overall? (A) No, only Westeros can (B) No, only Meereen can (C) Yes, both countries can consume more. (D) No, neither country benefits from trade.
Answer #1

1. Ans: A ) Meereen

2. Ans: B ) Westeros

3. Ans: A ) Meereen

Explanation:

Westeros 's opportunity cost of producing slaves = 3000 / 2000 = 1.5

Meereen's opportunity cost of producing slaves = 1000 / 3000 = 0.33

So it is cleared that Meereen's has comparative advantage in producing slaves due to lower opportunity cost .

4. Ans: D ) Westeros will produce 3,000 kg of wine, and Meereen will produce 3,000 slaves.

Explanation:

Westeros 's opportunity cost of producing slaves = 3000 / 2000 = 1.5

Meereen's opportunity cost of producing slaves = 1000 / 3000 = 0.33

Westeros 's opportunity cost of producing wine = 2000 / 3000 = 0.66

Meereen's opportunity cost of producing wine = 3000 / 1000 = 3

5. C ) Yes, both country can consume more.